### v1

@@ Abstract
Sequential recommender systems (SRS) could capture dynamic user preferences by modeling historical behaviors ordered in time. Despite effectiveness, focusing only on the collaborative signals from behaviors does not fully grasp user interests. It is also significant to model the semantic relatedness reflected in content features, e.g., images and text. Towards that end, in this paper, we aim to enhance the SRS tasks by effectively unifying collaborative signals and semantic relatedness together. Notably, we empirically point out that it is nontrivial to achieve such a goal due to semantic gap issues. Thus, we propose an end-to-end two-stream architecture for sequential recommendation, named TSSR, to learn user preferences from ID-based and content-based sequence. Specifically, we first present novel hierarchical contrasting module, including coarse user-grained and fine item-grained terms, to align the representations of inter-modality. Furthermore, we also design a two-stream architecture to learn the dependence of intra-modality sequence and the complex interactions of inter-modality sequence, which can yield more expressive capacity in understanding user interests. We conduct extensive experiments on five public datasets. The experimental results show that the TSSR could yield superior performance than competitive baselines. We also make our experimental codes publicly availablehttps://anonymous.4open.science/r/TSSR-2A27/.
@@ 1 Introduction
With the information overload, recommender systems [13, 32] have become one of the most useful information filtering engines for various online applications, such as e-commerce, advertising, and online videos. Among them, sequential recommender systems (SRS) [27, 5] have become ever-increasingly prevalent due to their strengths in capturing dynamic user interests relative to traditional collaborative filtering approaches. Formerly, many efforts [9] have been devoted to designing sophisticated models to learn precise user and item embeddings in the SRS problem, including early traditional methods, e.g., matrix factorization combined with the Markov chain [21], and recently various advanced neural models, including recurrent neural networks (RNNs) [9], convolutional neural networks (CNNs) [25], self-attentive models [12, 24].
@@
Though achieving success, the recommendation performance is unsatisfactory due to the sparse interactions of items. Luckily, various content features related to each item can be obtained in the SRS tasks [8], such as each item’s image appearance and textual descriptions. In general, visual contents play a significant role in influencing user decisions. Toward integrating content features to empower the SRS task, researchers have proposed some attempts in previous works [9, 34]. Nevertheless, we argue that these previous works only achieve limited performance gains in mining content features for sequential recommendation due to ignoring the semantic difference in embeddings of item IDs and content features. The main reason behind item ID is that corresponding learned user and item embeddings could represent the implicit collaborative signals among users, e.g., co-occurrence patterns. In contrast, content features typically contain very specific descriptions of an item. In real applications, it is very common to find similar items in terms of content features within a session. With the existence of semantic relatedness, these high-quality content feature can be utilized in performing recommendations even without leveraging item ID as shown in Table 1. Nevertheless, these two aspects of embeddings reflect the user preference from different views. Based on the above analysis, we set the goal of absorbing the capacity of collaborative signals and semantic relatedness together by treating item IDs and content features as two modalities. However, simply concatenating these features didn’t improve performance in our tests as shown in Table 1, likely due to the challenge of aligning both modalities in a shared latent space, which restricts the system’s effectiveness. To illustrate this, we used t-SNE to visualize item embeddings from ID-only and content-only recommendation methods, as displayed in Figure 1. The visualizations support our claim of a significant semantic gap between the embeddings of item IDs and content features. We assert that it’s essential to model not just the sequence dependencies within each modality but also the dependencies across modalities to fully grasp user interests. To that end, in this work, we propose an end-to-end two-stream architecture based on treating item IDs and content features as two modalities for sequential recommendations, refer to TSSR. The key idea behind the proposed model is a carefully designed hierarchical contrasting module for solving the issue of natural unaligned representations of item IDs and content features. Specifically, two levels of contrasting objectives, including coarse user-grained and fine item-grained terms, are proposed to align the representations of inter-modality. Besides, we introduce a two-stream encoder design, enabling not only sequence dependence on intra-modality but also item relationships of inter-modality to be well modeled, simultaneously. Due to comprehensively understanding user preferences via unifying collaborative signals and semantic signals, the TSSR naturally exhibits more accurate recommendations compared to previous approaches. We summarize the contributions as follows:
@@ 2.2 Empirical Studies
We posit that content feature embeddings are already effectively captured by existing visual and textual models [7], and it’s common to find similar items within a session. Contrary to expectations, as Table 1 reveals, content-only SRS performed comparably or even better than ID-only models. This suggests that the semantic relatedness within content feature sequences can significantly enhance recommendation quality. We believe this is reasonable because content embeddings encapsulate specific information, such as an item’s style, size, and color.
@@ 3.3 Aligning Unimodal Representations
To align two unimodal representations, we adopt a contrastive loss approach, proven effective in unsupervised learning breakthroughs [6, 18]. Item ID representations capture collaborative signals, while content features reflect semantic connections, both crucial for profiling user preferences. Therefore, we consider inter-modality embeddings as dual perspectives of the same user behavior, without relying on data augmentation. The calculation of the user-level contrastive loss $\mathcal{L}_{u}$ is detailed as follows,
@@ 3.4 Multimodal Interaction and Fusion
Building on the insights from recent studies [18], which highlight the advantages of leveraging multiple views of data for better object understanding, we aim to integrate multimodal information to enhance our comprehension of users’ interests. To this end, we process the modality-specific representations through two separate multimodal encoders running in parallel. The unique aspect of these multimodal encoders is the substitution of self-attention with cross-attention mechanisms. This approach involves using the representation from one modality as the query and the representation from the other as the key and value in the attention calculation. This design’s underlying principle is to enable each modality’s representation to engage with and incorporate information from the other modality. After $\rm L_{multi}$ layers multimodal aggregating, the sequence representations of IDs and contents are denoted by $\rm\overline{\textbf{F}}^{id},\overline{\textbf{F}}^{con}\in\mathbb{R}^{t\times d}$. To obtain the comprehensive item representations, we would merge the representation of two streams. Here, we devise a simple gate mechanism to learn the weight of each stream. After getting the output of sequence embedding, we project them into vectors by concatenation operation followed by a linear projection layer. Then, we compute the gate weight with the sigmoid function. Then each gating weight is attending to the corresponding stream’s output and is packed as the final feature vector. Formally, the whole fusion process can be described as,
$\begin{split}&\rm\textbf{S}=sigmoid(\textbf{W}[\overline{\textbf{F}}^{id},\overline{\textbf{F}}^{con}]+b),\\ &\rm\textbf{F}=(\textbf{S}\odot\overline{\textbf{F}}^{id})\oplus((\textbf{1}-\textbf{S})\odot\overline{\textbf{F}}^{con}),\end{split}$ (3)
where []delimited-[][$\cdot$][ ⋅ ] is the concatenation operation, and $\oplus$ denotes the element-wise add operation while $\odot$ indicates the element-wise multiplication calculation. $\textbf{W}\in\mathbb{R}^{2d\times d}$ is a projection matrix and $b$ indicates the bias term, and $\textbf{S}\in\mathbb{R}^{t\times d}$ denotes gate matrix.
@@ 3.5 Model Optimization
After obtaining the multimodal representation of whole sequence F, we next calculate each user’s preference score for the item $i$ in the step $(t+1)$ under the input context with the inner product by following [12]. For the next item recommendation prediction task, we employ the auto-regressive generative loss as optimization objectives, i.e., left-to-right supervision signals, whose effectiveness has been demonstrated in previous works [12]. In this work, we employ cross-entropy loss combined with such optimization objectives (denoted by $\mathcal{L}_{\rm ce}$) to help train the whole model. Lastly, the proposed TSSR architecture can be optimized in a multi-task manner as follows
@@
in which $\lambda_{1},\lambda_{2},\lambda_{3}$, denoting the strength of each objective.
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We conduct experiments on five public datasets:
  • H&M is a publicly available dataset released in Kaggle Challenge 2022111https://www.kaggle.com/c/h-and-m-personalized-fashion-recommendations. We take the historical purchase behaviors into account for constructing sequence recommendation datasets, in which each item’s image appearance is regarded as the corresponding content features.
  • Yelp is a public dataset  [^2] https://www.yelp.com/dataset, where we obtain each user’s sequential behaviors by ranking their rated items in time order. The image appearance is treated as content features of items.
  • Phone and Toy are two datasets selected from Amazon review data [^3] http://deepyeti.ucsd.edu/jianmo/amazon/index.html, consists of “Cell Phones and Accessories” (Phone), “Toys and Games” (Toy), respectively. And we construct sequential behaviors by using the user’s rating histories and regard the image of each item as the contents.
  • MIND is an English dataset for monolingual news recommendation [^4] https://msnews.github.io/. For each user, their history click logs are used to construct sequential behaviors, whose title textual descriptions are treated as contents.
We set the fixed sequence length of 10 for four datasets (including H&M Yelp, Phone, Toy) while setting MIND as $20$. Besides, we filter items and users with fewer than five interaction records. And we randomly pick $100$ thousand users from the H&M dataset in experiments. The statics of these datasets after processing are summarized in Table 2.
@@ 4.1.2 Compared Methods
To verify the effectiveness of the TSSR, we choose representative methods as baselines, including traditional methods and deep sequential models. In particular, we regard some hybrid methods as baselines, in which both content and interaction records are simultaneously modeled. Note that we do not compare recent graph-based methods [31] which are mainly studied in the setting of session-based recommendation while our work focuses on sequential recommendation setting [27]. Next, we detail the compared baselines,
@@
  • MFN4Rec [28] treat the content feature as auxiliary side information by employing a multi-view gated memory network for the sequential recommendation.
  • SASRec-ID [12] is a powerful SRS method based on self-attentive-based architectures, i.e., a transformer encoder to capture long-range dependence of sequential behaviors.
  • SASRec-Content [12] is a variant of SASRec, in which we directly regard the content feature representation as sequence input. This baseline is an ID-agnostic baseline.
Noted that it preserves the same input between several content-enriched recommendation models and our proposed TSSR.
4.1.3 Evaluation Metrics.
4.1.4 Hyper-parameter Settings
4.2 Analysis of Experimental Results
@@ 4.2.1 Recommendation Performance Evaluation
The results in Table 3 show that TSSR outperforms the competitive baselines, confirming the effectiveness of our approach in jointly modeling item ID and content feature sequences. Other content-rich SRS models fall short due to a lack of representation alignment or effective modeling of complex relations within or across modalities, resulting in sub-optimal performance. In contrast, TSSR exceeds previous methods by thoroughly analyzing historical user behavior. Additionally, we observe that TSSR particularly excels in the clothing domain, likely because visual elements are more influential in clothing selection than in other areas. However, improvements in scenarios with textual content are more modest, possibly because only news titles are considered as content. This suggests that developing methods for extracting textual representations is a promising avenue for future research.
@@ 4.2.2 Study of Representation Alignment
In this study, we delve deeper into the representation module of the full TSSR model by conducting ablation experiments. We test variants of the model where we remove the coarse user-level, fine item-level, or all alignment operations, respectively. We evaluate these variants on the H&M dataset, focusing on two key metrics: Recall@20 and NDCG@20. Additionally, we assess the quality of the learned representations in terms of alignment and uniformity[29], which measure the closeness of paired instance embeddings and the even distribution of embeddings, respectively. It’s important to note that we normalize all representations before evaluation. Our findings, summarized in Figure 3, show that the full representation alignment optimization indeed enhances the TSSR model. Interestingly, we observe that better alignment and uniformity in representations do not necessarily lead to improved recommendation performance and may even hinder it. We attribute these results to two main factors: first, the similarity between item-grained contrasting and generative auto-regressive objectives; second, the intrinsic nature of item recommendation as a ranking rather than a classification problem  [4]. To further intuitively show how well the item embeddings are aligned by our hierarchical contrasting module, we show the fusion performance of learned item embeddings by performing t-SNE over-extracted embeddings. As shown in Figure 4, the left figure denotes the t-SNE visualization comparison across embeddings of item IDs and content features while the right figure indicates the visualization of aligned embeddings in the TSSR model. Such results further demonstrate the effectiveness of our TSSR in aligning representations of item IDs and content features.
@@
To solve the semantic gap issue, we carefully develop the hierarchical contrasting module, in which the number of negative sampling plays a vital role in the final performance. Due to the mini-batch sharing strategy in our work, we hence turn to empirically evaluate how would batch size affects performance. For clarity and saving space, we only report the performance of TSSR trained in the Phone dataset as shown in Figure 5. Note that here we not only evaluate the recommendation results but also examine the representation quality of learned embedding with alignment and uniformity. From the experimental results, we find that: 1) our proposed TSSR can benefit from a larger batch size, and 2) the batch size can also greatly influence the value of uniformity. We hold such an experimental phenomenon is reasonable since that larger batch size can provide more negative examples for hierarchical contrasting optimization. Hence, the representation quality of item representations is naturally improved with the increasing of negative examples.
@@ 4.2.5 Limitations Analysis
Although our work can effectively organize item IDs and content features in a unified framework, we also notice two specific limitations in our work. One key limitation is the high training cost of TSSR because of the end-to-end learning paradigm. For example, TSSR typically requires around 120G GPU memory in the H&M dataset. In contrast, around 2G GPU memory is enough for ID-only baselines (i.e., SASRec-ID). Luckily, we notice some adapter-based parameter-efficient learning techniques have been extensively studied, in which a very small proportion of trainable model parameters (e.g., only $3\%$ in the classic Houlsby adapter [11]) is able to achieve comparable performances. A second limitation is that we only examine TSSR with one type of content feature for each item. We are aware of aggregating multiple categories of content features into new feature sequences might benefit the final sequential recommendation performance. In future work, we plan to evaluate the effectiveness of our work by collecting multiple types of modality features of items.

### v2

@@ Abstract
Sequential recommender systems (SRS) could capture dynamic user preferences by modeling historical behaviors ordered in time. Despite effectiveness, focusing only on the collaborative signals from behaviors does not fully grasp user interests. It is also significant to model the semantic relatedness reflected in content features, e.g., images and text. Towards that end, in this paper, we aim to enhance the SRS tasks by effectively unifying collaborative signals and semantic relatedness together. Notably, we empirically point out that it is nontrivial to achieve such a goal due to semantic gap issues. Thus, we propose an end-to-end two-stream architecture for sequential recommendation, named TSSR, to learn user preferences from ID-based and content-based sequence. Specifically, we first present novel hierarchical contrasting module, including coarse user-grained and fine item-grained terms, to align the representations of inter-modality. Furthermore, we also design a two-stream architecture to learn the dependence of intra-modality sequence and the complex interactions of inter-modality sequence, which can yield more expressive capacity in understanding user interests. We conduct extensive experiments on five public datasets. The experimental results show that the TSSR could yield superior performance than competitive baselines. We also make our experimental codes publicly available at https://github.com/Mingyue-Cheng/TSSR.
@@ 1 Introduction
With the information overload, recommender systems [14, 31, 34] have become one of the most useful information filtering engines for various online applications, such as e-commerce, advertising, and online videos. Among them, sequential recommender systems (SRS) [27, 7, 26] have become ever-increasingly prevalent due to their strengths in capturing dynamic user interests relative to traditional collaborative filtering approaches. Formerly, many efforts [10] have been devoted to designing sophisticated models to learn precise user and item embeddings in the SRS problem, including early traditional methods, e.g., matrix factorization combined with the Markov chain [21], and recently various advanced neural models, including recurrent neural networks (RNNs), convolutional neural networks (CNNs), self-attentive models [13].
@@
Though achieving success, the recommendation performance is unsatisfactory due to the sparse interactions of items. Luckily, various content features related to each item can be obtained in the SRS tasks [9], such as each item’s image appearance and textual descriptions. In general, visual contents play a significant role in influencing user decisions. Toward integrating content features to empower the SRS task, researchers have proposed some attempts in previous works [10, 33]. Nevertheless, we argue that these previous works only achieve limited performance gains in mining content features for sequential recommendation due to ignoring the semantic difference in embeddings of item IDs and content features. The main reason behind item ID is that corresponding learned user and item embeddings could represent the implicit collaborative signals among users, e.g., co-occurrence patterns. In contrast, content features typically contain very specific descriptions of an item. In real applications, it is very common to find similar items in terms of content features within a session. With the existence of semantic relatedness, these high-quality content feature can be utilized in performing recommendations even without leveraging item ID as shown in Table 1. Nevertheless, these two aspects of embeddings reflect the user preference from different views.
Based on the above analysis, we set the goal of absorbing the capacity of collaborative signals and semantic relatedness together by treating item IDs and content features as two modalities. However, simply concatenating these features didn’t improve performance in our tests as shown in Table 1, likely due to the challenge of aligning both modalities in a shared latent space, which restricts the system’s effectiveness. To illustrate this, we used t-SNE to visualize item embeddings from ID-only and content-only recommendation methods, as displayed in Figure 1. The visualizations support our claim of a significant semantic gap between the embeddings of item IDs and content features. We assert that it’s essential to model not just the sequence dependencies within each modality but also the dependencies across modalities to fully grasp user interests.
To that end, in this work, we propose an end-to-end two-stream architecture based on treating item IDs and content features as two modalities for sequential recommendations, refer to TSSR. The key idea behind the proposed model is a carefully designed hierarchical contrasting module for solving the issue of natural unaligned representations of item IDs and content features. Specifically, two levels of contrasting objectives, including coarse user-grained and fine item-grained terms, are proposed to align the representations of inter-modality. Besides, we introduce a two-stream encoder design, enabling not only sequence dependence on intra-modality but also item relationships of inter-modality to be well modeled, simultaneously. Due to comprehensively understanding user preferences via unifying collaborative signals and semantic signals, the TSSR naturally exhibits more accurate recommendations compared to previous approaches. We summarize the contributions as follows:
@@ 2.2 Empirical Studies
We posit that content feature embeddings are already effectively captured by existing visual and textual models [8], and it is common to find similar items within a session. Contrary to expectations, as Table 1 reveals, content-only SRS performed comparably or even better than ID-only models. This suggests that the semantic relatedness within content feature sequences can significantly enhance recommendation quality. We believe this is reasonable because content embeddings encapsulate specific information, such as an item’s style, size, and color.
@@ 3.3 Aligning Unimodal Representations
To align two unimodal representations, we adopt a contrastive loss approach, proven effective in unsupervised learning breakthroughs [18]. Item ID representations capture collaborative signals, while content features reflect semantic connections, both crucial for profiling user preferences. Therefore, we consider inter-modality embeddings as dual perspectives of the same user behavior, without relying on data augmentation. The calculation of the user-level contrastive loss $\mathcal{L}_{u}$ is detailed as follows,
@@ 3.4 Multimodal Interaction and Fusion
Building on the insights from recent studies [18], which highlight the advantages of leveraging multiple views of data for better object understanding, we aim to integrate multimodal information to enhance our comprehension of users’ interests. To this end, we process the modality-specific representations through two separate multimodal encoders running in parallel. The unique aspect of these multimodal encoders is the substitution of self-attention with cross-attention mechanisms. This approach involves using the representation from one modality as the query and the representation from the other as the key and value in the attention calculation. This design’s underlying principle is to enable each modality’s representation to engage with and incorporate information from the other modality. After $\rm L_{multi}$ layers multimodal aggregating, the sequence representations of IDs and contents are denoted by $\rm\overline{\textbf{F}}^{id},\overline{\textbf{F}}^{con}\in\mathbb{R}^{t\times d}$.
To obtain the comprehensive item representations, we would merge the representation of two streams. Here, we devise a simple gate mechanism to learn the weight of each stream. After getting the output of sequence embedding, we project them into vectors by concatenation operation followed by a linear projection layer. Then, we compute the gate weight with the sigmoid function. Then each gating weight is attending to the corresponding stream’s output and is packed as the final feature vector. Formally, the whole fusion process can be described as,
$\begin{split}&\rm\textbf{S}=sigmoid(\textbf{W}[\overline{\textbf{F}}^{id},\overline{\textbf{F}}^{con}]+b),\\ &\rm\textbf{F}=(\textbf{S}\odot\overline{\textbf{F}}^{id})+((\textbf{1}-\textbf{S})\odot\overline{\textbf{F}}^{con}),\end{split}$ (3)
where []delimited-[][$\cdot$][ ⋅ ] is the concatenation operation, and $\odot$ indicates the element-wise multiplication calculation. $\textbf{W}\in\mathbb{R}^{2d\times d}$ is a projection matrix and $b$ indicates the bias term, and $\textbf{S}\in\mathbb{R}^{t\times d}$ denotes gate matrix.
@@ 3.5 Model Optimization
After obtaining the multimodal representation of whole sequence F, we next calculate each user’s preference score for the item $i$ in the step $(t+1)$ under the input context with the inner product by following [13, 2]. For the next item recommendation prediction task, we employ the auto-regressive generative loss as optimization objectives, i.e., left-to-right supervision signals, whose effectiveness has been demonstrated in previous works [1]. In this work, we employ cross-entropy loss combined with such optimization objectives (denoted by $\mathcal{L}_{\rm ce}$) to help train the whole model. Lastly, the proposed TSSR architecture can be optimized in a multi-task manner with the trade-off parameters $\lambda_{1},\lambda_{2},\lambda_{3}$ for each objective as follows:
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We conduct experiments on five public datasets:
  • H&M is a publicly available dataset released in Kaggle Challenge 2022https://www.kaggle.com/c/h-and-m-personalized-fashion-recommendations. We take the historical purchase behaviors into account, in which each item’s image appearance is regarded as the corresponding content features. And we randomly pick $100$ thousand users in experiments.
  • Yelp is a public dataset https://www.yelp.com/dataset, where we obtain each user’s sequential behaviors by ranking their rated items in time order. The image appearance is treated as content features of items.
  • Phone and Toy are two datasets selected from Amazon review datahttp://deepyeti.ucsd.edu/jianmo/amazon/index.html, consists of “Cell Phones and Accessories” (Phone), “Toys and Games” (Toy), respectively. And we construct sequential behaviors by using the user’s rating histories and regard the image of each item as the contents.
  • MIND is an English dataset for monolingual news recommendationhttps://msnews.github.io/. For each user, their history click logs are used to construct sequential behaviors, whose title textual descriptions are treated as contents.
@@ 4.1.1 Datasets
We set the fixed sequence length of 20 for MIND while setting other datasets as $10$. Besides, we filter items and users with fewer than five interaction records. The statics of these datasets after processing are summarized in Table 2.
@@ 4.1.2 Compared Methods
To verify the effectiveness of the TSSR, we choose representative methods as baselines, including traditional methods and deep sequential models. In particular, we regard some hybrid methods as baselines, in which both content and interaction records are simultaneously modeled. Note that we do not compare recent graph-based methods which are mainly studied for session-based recommendation. Next, we detail the compared baselines,
@@
  • MFN4Rec [28] treat the content feature as the side information by employing a multi-view gated memory network for sequential recommendation.
  • SASRec-ID [13] is a powerful SRS method based on self-attentive-based architectures to capture long-range dependence of sequential behaviors.
  • SASRec-Content [13] is a variant of SASRec, in which we directly regard the content feature representation as sequence input.
@@
4.1.3 Evaluation Metrics.
4.1.4 Hyper-parameter Settings
4.2 Analysis of Experimental Results
@@ 4.2.1 Recommendation Performance Evaluation
The results in Table 3 show that TSSR outperforms the competitive baselines, confirming the effectiveness of our approach in jointly modeling item ID and content feature sequences. Other content-rich SRS models fall short due to a lack of representation alignment or effective modeling of complex relations within or across modalities, resulting in sub-optimal performance. Additionally, we observe that TSSR particularly excels in the clothing domain, likely because visual elements are more influential in clothing selection than in other areas. However, improvements in scenarios with textual content are more modest, possibly because only news titles are considered as content. This suggests that developing methods for extracting textual representations is a promising avenue for future research.
@@ 4.2.2 Study of Representation Alignment
In this study, we delve deeper into the representation module of the full TSSR model by conducting ablation experiments. We test variants of the model where we remove the coarse user-level, fine item-level, or all alignment operations, respectively. We evaluate these variants on the H&M dataset. Additionally, we assess the quality of the learned representations in terms of alignment and uniformity [29], which measure the closeness of paired instance embeddings and the even distribution of embeddings, respectively. It’s important to note that we normalize all representations before evaluation. Our findings, summarized in Figure 3, show that the full representation alignment optimization indeed enhances the TSSR model. Interestingly, we observe that better alignment and uniformity in representations do not necessarily lead to improved recommendation performance and may even hinder it. We attribute these results to two main factors: first, the similarity between item-grained contrasting and generative auto-regressive objectives; second, the intrinsic nature of item recommendation as a ranking rather than a classification problem  [6]. To further intuitively show how well the item embeddings are aligned by our hierarchical contrasting module, we show the fusion performance of learned item embeddings by performing t-SNE over-extracted embeddings. As shown in Figure 4, the left figure denotes the t-SNE visualization comparison across embeddings of item IDs and content features while the right figure indicates the visualization of aligned embeddings in the TSSR model. Such results further demonstrate the effectiveness of our TSSR in aligning representations of item IDs and content features.
@@
To solve the semantic gap issue, we carefully develop the hierarchical contrasting module, in which the number of negative sampling plays a vital role in the final performance. Due to the mini-batch sharing strategy in our work, we hence turn to empirically evaluate how would batch size affects performance. For clarity and saving space, we only report the performance of TSSR trained in the Phone dataset as shown in Figure 5. Note that we also examine the representation quality of learned embedding with alignment and uniformity. From the experimental results, we find that: 1) our proposed TSSR can benefit from a larger batch size, and 2) the batch size can also greatly influence the value of uniformity. We hold such an experimental phenomenon is reasonable since that larger batch size can provide more negative examples for hierarchical contrasting optimization. Hence, the representation quality of item representations is naturally improved with the increasing of negative examples.
4.2.4 Result Analysis w.r.t. the Amount of Training Data
@@ 4.2.5 Limitations Analysis
Although our work can effectively organize item IDs and content features in a unified framework, we also notice the key limitation about high training cost of TSSR because of the end-to-end learning paradigm. For example, TSSR typically requires around 120G GPU memory in the H&M dataset. In contrast, around 2G GPU memory is enough for ID-only baselines (i.e., SASRec-ID). Luckily, we notice some adapter-based parameter-efficient learning techniques have been extensively studied, in which a very small proportion of trainable model parameters (e.g., only $3\%$ in the classic Houlsby adapter [12]) is able to achieve comparable performances.